OpenStudy (anonymous):

Help please! Write an equation of the line, in point-slope form, that passes through the given points: (-2,15), (9,-18).

OpenStudy (anonymous):

First find slope: Slope formula:\[m=\frac{ (y2-y1) }{ (x2-x1)}\] (-2,15), (9,-18). x1,y1 x2,y2

OpenStudy (anonymous):

-33/7 for my slope? @Munch.

OpenStudy (anonymous):

-18-15/9-(-2) -33/11

OpenStudy (anonymous):

-3 is the slope.

OpenStudy (anonymous):

Now to find the equation use this formula \[y-y1=m(x-x1)\]

OpenStudy (anonymous):

m=slope that is 3 and you know other values.

OpenStudy (anonymous):

y - 15 = -3(x - 2)?

OpenStudy (anonymous):

x+2 because double negative is postive.